OpenStudy (anonymous):

A rubber ball is released from a height of 180cm above the surface of the floor. The ball bounces back 4/5 of its previous height each time. Find the total distance traveled by the ball before it stops.

OpenStudy (anonymous):

180*4/5=144 144/(1-(4/5))=144/(1/5)=144*5=720 720*2=1440 1440+180=1620